# UAV-UGV Cooperative Trajectory Optimization and Task Allocation for Medical Rescue Tasks in Post-Disaster Environments

### v1

@@ II-E Problem Scope and Formal Summary
III Methodology
III-A Task Allocation via Enhanced Genetic Algorithm (EGA)
@@
The optimization objective is formulated as:
@@
$\bigcup_{i=1}^{N}\mathcal{T}_{i}=\mathcal{T},\quad\mathcal{T}_{i}\cap\mathcal{T}_{j}=\emptyset\quad\forall i\neq j,i=1,$ (6)
@@
To solve this optimization problem, an enhanced GA is utilized. Each individual in the GA population represents a candidate task allocation encoded as a sequence of task indices grouped by assigned vehicles. The fitness of an individual is evaluated according to the total path cost defined by the above objective function.
A selection operator based on tournament selection is employed to favor individuals with lower total cost, while maintaining diversity. The crossover operator partially exchanges assigned tasks between two parents. After crossover, a repair mechanism is applied to ensure all tasks remain uniquely assigned. Mutation is performed by randomly swapping tasks between different vehicles or rearranging local task orders. To adaptively balance exploration and exploitation, the mutation probability $\mu$ is adjusted over generations:
@@
where $\mu_{0}$ is the initial mutation probability, $g$ is the current generation number, $G$ is the total number of generations, and $\alpha$ controls the decay rate. Furthermore, an elite preservation mechanism ensures that the top $E$of individuals with the lowest cost are directly carried over to the next generation, preventing the loss of high-quality solutions. The overall GA procedure iteratively improves the population until a stopping criterion is met, such as reaching a maximum number of generations or achieving convergence in the fitness values. Through these enhancements, the GA effectively navigates the large and complex search space of multi-vehicle task allocation under post-disaster conditions, producing high-quality allocations that significantly reduce mission time and energy consumption.
III-B Trajectory Generation via Informed-RRT*
@@
III-C Trajectory Optimization with CMA-ES
@@
$\min_{\pi_{i}}\quad C_{i}(\pi_{i})=\sum_{j=0}^{n_{i}-1}d(P_{\pi_{i}(j)},P_{\pi_{i}(j+1)}),$ (11)
where $\pi_{i}(j)$ denotes the index of the $j$-th task in the visiting sequence, and $d(\cdot,\cdot)$ is the Euclidean distance between two consecutive task locations.
@@
III-D Integrated Cooperative Planning Framework
The integrated planning framework commences by modeling the post-disaster environment, defining obstacle regions $\mathcal{O}$, task points $\mathcal{T}$, and the base station B. Subsequently, an EGA is employed to allocate medical rescue tasks $\mathcal{T}$ among the heterogeneous fleet of UAVs $\mathcal{U}$ and UGVs $\mathcal{V}$, aiming to optimize overall mission efficiency. Following task allocation, the Informed-RRT* algorithm generates initial collision-free and constraint-compliant trajectories $P_{i}(t)$ for each vehicle $i$ to visit its assigned tasks. The CMA-ES is then utilized to optimize the task visiting sequence $\pi_{i}$ and refine the trajectory for each vehicle, minimizing the total mission completion time $T_{\text{total}}$ and the sum of path lengths or energy consumption $\sum C(P_{i})$. Finally, the complete plan is evaluated to ensure adherence to all constraints, including safety distances $d_{\text{safe}}$, dynamics, energy limits $E_{i}$, and full task coverage.
@@ III-D Integrated Cooperative Planning Framework
This integrated approach offers several advantages for the target application. By separating concerns—global allocation (EGA), feasible pathfinding (Informed-RRT*), and local sequence optimization (CMA-ES)—it decomposes the complex overall problem into manageable subproblems. It systematically leverages the strengths of different optimization paradigms suited for different aspects of the problem (combinatorial allocation, sampling-based path planning, continuous optimization for sequencing). This structured integration facilitates the generation of globally coordinated, locally optimized, feasible, and efficient plans for heterogeneous multi-robot teams operating in challenging post-disaster environments, ultimately aiming to improve the timeliness and effectiveness of medical rescue operations.
IV Simulation Results
IV-A Simulation Setup
To validate the effectiveness and robustness of the proposed cooperative UAV-UGV trajectory optimization and task allocation framework, extensive simulations are conducted in a realistic post-disaster environment.
@@ IV-A Simulation Setup
The primary metrics used to evaluate the performance of the proposed framework include: 1.Total Mission Completion Time ($T_{\text{total}}$): Time elapsed until the last vehicle completes its assigned tasks and returns to the base. 2.Total Path Length ($\sum L_{i}$): Sum of the trajectory lengths for all vehicles, reflecting overall travel distance and related energy consumption. 3.Energy Consumption Estimation ($\sum C(P_{i})$): Calculated based on path lengths and vehicle-specific energy models. Feasibility, including collision avoidance and constraint satisfaction, is a prerequisite for all generated plans.
IV-B Results and Analysis
Comparative experiments were designed to evaluate the EGA-based task allocation against baseline methods such as random assignment, K-Means clustering and GA Optimization Allocation. Expected results, visualized using bar charts Fig. 1, will demonstrate that the EGA significantly reduces the total path length and/or mission completion time compared to these simpler strategies, highlighting its effectiveness in finding efficient assignments under complex constraints.
Figure 1: task assignment comparison
@@ IV-B Results and Analysis
Figure 2: convergence of genetic algorithm for task allocation
@@
Fig. 3 and 4 provide compelling evidence for the value of the CMA-ES trajectory optimization stage. By intelligently reordering the sequence in which tasks are visited, CMA-ES effectively shortens the required travel distances and streamlines the routes, thereby significantly improving the overall efficiency and likely reducing the energy consumption of the planned missions compared to using only the initial feasible paths from Informed-RRT*.
Figure 3: path optimization comparison
Figure 4: path optimization distance comparison
@@
TABLE I: Extensibility Analysis
@@
V Conclusion
This paper proposed a cooperative trajectory optimization and task allocation framework for heterogeneous UAV-UGV teams operating in post-disaster medical rescue scenarios. By integrating a genetic algorithm for efficient task distribution, Informed-RRT* for feasible path generation, and CMA-ES for trajectory sequence optimization, the framework effectively addresses the challenges of dynamic environments with complex obstacle distributions. Simulations conducted in a realistic disaster scenario demonstrate that the proposed method significantly reduces total mission completion time and overall path length compared to baseline strategies. Through efficient collaboration between UAVs and UGVs, the system enhances operational flexibility, improves task execution efficiency, and ensures feasible and continuous trajectories suitable for real-world deployment. The results validate the effectiveness and robustness of the approach in achieving efficient rescue operations under constrained and hazardous conditions.
VI Code Availability

### v2

@@ 3.5 Problem Scope and Formal Summary
4 Methodology
4.1 Task Allocation via Enhanced Genetic Algorithm (EGA)
@@
GAs are a class of evolutionary algorithms inspired by natural selection and genetics. They are particularly well-suited for complex combinatorial optimization problems like ours, where the search space is vast and discontinuous. A GA works by maintaining a population of candidate solutions (individuals), which evolve over successive generations through processes of selection, crossover, and mutation. High-fitness solutions (those with lower mission cost) are more likely to be selected for reproduction, leading the population to converge towards optimal or near-optimal solutions.
To solve this optimization problem, we propose an EGA that builds upon the standard GA framework with several key improvements to enhance convergence speed and solution quality. The optimization objective is formulated as:
@@
$\bigcup_{i=1}^{N}T_{i}=T,\quad T_{i}\cap T_{j}=\emptyset\quad\forall i\neq j,\quad i,j\in\{1,\dots,N\},$ (6)
@@
To illustrate this encoding, consider a scenario with 3 vehicles (2 UAVs: $U_{1}$, $U_{2}$; 1 UGV: $G_{1}$) and 7 tasks ($T_{1}$ to $T_{7}$). A possible individual (chromosome) could be represented as: $\text{Chromosome: }[U_{1}:T_{1},T_{3},T_{5}\mid U_{2}:T_{2},T_{6}\mid G_{1}:T_{4},T_{7}]$.
In this example, the chromosome is a single sequence divided by vehicle identifiers. $U_{1}$ is assigned the task sequence $T_{1}\rightarrow T_{3}\rightarrow T_{5}$, $U_{2}$ is assigned $T_{2}\rightarrow T_{6}$, and $G_{1}$ is assigned $T_{4}\rightarrow T_{7}$. This encoding explicitly captures both the assignment of tasks to vehicles and the sequence in which those tasks are visited.
The EGA is designed as follows: Each individual in the GA population is represented as a sequence of task indices, grouped by the vehicle to which they are assigned. The fitness of an individual is evaluated according to the total path cost defined by the objective function (5). A tournament selection operator is employed to favor individuals with lower total cost. A partially matched crossover (PMX) operator is used to exchange assigned tasks between two parent individuals, followed by a repair mechanism to ensure feasibility. Mutation is performed by randomly swapping tasks between different vehicles or rearranging the local task order within a vehicle’s sequence.
The enhancements in our EGA are two-fold:
1. Adaptive Mutation Probability: To balance exploration in early generations with exploitation in later generations, the mutation probability $\mu$ is dynamically adjusted over time:
@@
where $\mu_{0}$ is the initial mutation probability, $g$ is the current generation number, $G$ is the total number of generations, and $\alpha$ controls the decay rate. This allows for aggressive exploration initially and fine-tuning later.
2. Elite Preservation: An elite preservation mechanism ensures that the top $E\%$ of individuals with the lowest cost are directly carried over to the next generation. This prevents the loss of high-quality solutions and accelerates convergence.
The overall EGA procedure iteratively improves the population until a stopping criterion is met. The pseudo-code for the proposed EGA is presented in Algorithm 1.
Algorithm 1 Enhanced Genetic Algorithm for Task Allocation
0:  Task set $\mathcal{T}$, Vehicle set $\mathcal{V}$, Population size $P$, Max generations $G$, Initial mutation rate $\mu_{0}$, Elite ratio $E$
0:  Optimal task assignment $X_{\text{best}}$
  Initialize population $P_{\text{pop}}$ with $P$ random individuals
  Evaluate fitness for all individuals in $P_{\text{pop}}$
  $X_{\text{best}}\leftarrow$ best individual from $P_{\text{pop}}$
  for $g=1$ to $G$ do
   Select parents from $P_{\text{pop}}$ using tournament selection
   Perform PMX crossover to generate offspring
   Apply repair mechanism to ensure task uniqueness
   Calculate mutation probability $\mu(g)$ using Eq. (7)
   Perform mutation on offspring with probability $\mu(g)$
   Evaluate fitness for new offspring
   Form new population: Combine new offspring with $E\%$ elites from $P_{\text{pop}}$
   $P_{\text{pop}}\leftarrow$ new population
   Update $X_{\text{best}}$ if a better individual is found
  end for
  return  $X_{\text{best}}$
4.2 Trajectory Generation via Informed-RRT*
@@
The core of the Informed-RRT* algorithm lies in its dynamically defined sampling domain. Once an initial feasible path with cost $c_{\text{best}}$ is found, the sampling is restricted to an ellipsoidal region $X_{\text{informed}}$, mathematically defined as:
$X_{\text{informed}}=\{\mathbf{x}\in\mathbb{R}^{2}\mid\|\mathbf{x}-\mathbf{x}_{\text{start}}\|+\|\mathbf{x}-\mathbf{x}_{\text{goal}}\|\leq c_{\text{best}}\}$ (11)
where $\mathbf{x}_{\text{start}}$ and $\mathbf{x}_{\text{goal}}$ are the start and goal positions, respectively. This constraint ensures that only states potentially leading to a path with a cost lower than $c_{\text{best}}$ are sampled, which is the key mechanism for its accelerated convergence.
The use of Informed-RRT* provides significant advantages over the standard RRT*. By focusing the sampling effort within the informed set $X_{\text{informed}}$, the algorithm drastically reduces the number of samples wasted in regions of the state space that cannot yield an improved solution. This leads to a much faster convergence rate towards a near-optimal path. Furthermore, like its predecessor, Informed-RRT* retains the theoretical property of asymptotic optimality under certain conditions, guaranteeing that the solution cost will converge to the global optimum with probability one given sufficient time, providing a strong theoretical foundation for the trajectory generation phase.
@@
4.3 Trajectory Optimization with CMA-ES
@@
$\min_{\pi_{i}}C_{i}(\pi_{i})=\sum_{j=0}^{n_{i}-1}d\big{(}P_{\pi_{i}(j)},P_{\pi_{i}(j+1)}\big{)},$ (12)
where $n_{i}$ denotes the number of tasks assigned to vehicle $V_{i}$, $\pi_{i}$ is a permutation of the task indices that specifies the visiting order, with $\pi_{i}(j)$ indicating the index of the $j$-th task visited by $V_{i}$. The 2D coordinate of the $\pi_{i}(j)$-th task is denoted by $P_{\pi_{i}(j)}$, and we define $P_{\pi_{i}(0)}=P_{\pi_{i}(n_{i})}=B$, where $B$ is the base location (depot), ensuring that each vehicle starts and ends its route at the depot. The function $d(\cdot,\cdot)$ computes the Euclidean distance between two points in the plane. Finally, $C_{i}(\pi_{i})$ represents the total path cost for vehicle $V_{i}$, defined as the sum of the travel distances along the route $\pi_{i}$, i.e., the total distance traveled by $V_{i}$ when executing the task sequence $\pi_{i}$. This cost is the objective function minimized by the CMA-ES algorithm.
@@
CMA-ES is particularly well-suited for the trajectory sequence optimization problem due to its powerful adaptive mechanism. Unlike algorithms with fixed search strategies, CMA-ES learns the structure of the optimization landscape by continuously adapting its covariance matrix $C$. This allows it to efficiently identify and exploit correlations between decision variables (i.e., task positions in the sequence), effectively performing a principled local search in promising directions. Its ability to handle non-linear, non-convex, and potentially noisy objective functions makes it robust for refining paths in complex environments. The continuous relaxation approach bridges the gap between the discrete nature of the TSP-like problem and the continuous optimization capability of CMA-ES, enabling a sophisticated search that can escape local minima more effectively than simpler heuristics, ultimately leading to smoother and more efficient final trajectories.
@@
4.4 Integrated Cooperative Planning Framework
The integrated planning framework commences by modeling the post-disaster environment, defining obstacle regions $O$, task points $T$, and the base station $B$. Subsequently, an EGA is employed to allocate medical rescue tasks $T$ among the heterogeneous fleet of UAVs $U$ and UGVs $G$, aiming to optimize overall mission efficiency. Following task allocation, the Informed-RRT* algorithm generates initial collision-free and constraint-compliant trajectories $P_{i}(t)$ for each vehicle $i$ to visit its assigned tasks. CMA-ES is then utilized to optimize the task visiting sequence $\pi_{i}$ and refine the trajectory for each vehicle, minimizing the total mission completion time $T_{\text{total}}$ and the sum of path costs $\sum C_{i}(\pi_{i})$ (which directly relates to energy consumption $\sum E_{i}$). Finally, the complete plan is evaluated to ensure adherence to all constraints, including safety distances $d_{\text{safe}}$, dynamics, energy limits $E_{i}$, and full task coverage.Fig. 2 shows the integration framework diagram, showing the connections between the modules.
Figure 2: Integrated framework diagram
@@ 4.4 Integrated Cooperative Planning Framework
This integrated approach offers several advantages for the target application. By separating concerns—global allocation, feasible pathfinding, and local sequence optimization decomposes the complex overall problem into manageable subproblems. It systematically leverages the strengths of different optimization paradigms suited for different aspects of the problem . This structured integration facilitates the generation of globally coordinated, locally optimized, feasible, and efficient plans for heterogeneous multi-robot teams operating in challenging post-disaster environments, ultimately aiming to improve the timeliness and effectiveness of medical rescue operations.
5 Simulation Results
5.1 Simulation Setup
To validate the framework, simulations are conducted in a realistic post-disaster environment. The heterogeneous fleet includes UAVs and UGVs with distinct operational limits. All vehicles must adhere to their dynamic constraints, energy limits ($E_{i}$), and maintain a minimum safety distance $d_{\text{safe}}$ from obstacles and other vehicles. The energy constraint for each vehicle is enforced as:
$C_{i}(\pi_{i})\leq E_{i},$ (15)
where $C_{i}(\pi_{i})$ is the total path cost (in distance units) and $E_{i}$ is the maximum allowable cost (converted from energy budget via a vehicle-specific efficiency model). This ensures that longer paths, which consume more energy, are penalized during planning.
@@ 5.1 Simulation Setup
The primary metrics used to evaluate the performance of the proposed framework include: 1.Total Mission Completion Time ($T_{\text{total}}$): Time until the last vehicle returns to base. 2.Total Path Length ($\sum L_{i}$): Sum of all vehicle trajectories. For clarity, $L_{i}\equiv C_{i}(\pi_{i})$, so $\sum L_{i}=\sum C_{i}(\pi_{i})$. 3.Energy Consumption Estimation ($\sum E_{i}$): Estimated as $\sum E_{i}=\sum\eta_{i}\cdot C_{i}(\pi_{i})$, where $\eta_{i}$ is the energy consumption rate (e.g., J/m) for vehicle $V_{i}$. Thus, energy is directly proportional to the optimized path cost $C_{i}(\pi_{i})$.
5.2 Results and Analysis
Figure 3: Comparison of total path length generated by different task allocation strategies. The proposed Enhanced Genetic Algorithm (EGA) achieves the shortest total path length, demonstrating its superior efficiency in optimizing the overall mission route compared to Random Assignment, K-Means clustering and standard GA.
Comparative experiments were designed to evaluate the EGA-based task allocation against baseline methods such as random assignment, K-Means clustering, improved GA and standard GA optimization allocation. Expected results, visualized using bar charts Fig. 3, will demonstrate that the EGA significantly reduces the total path length and/or mission completion time compared to these simpler strategies, highlighting its effectiveness in finding efficient assignments under complex constraints.
@@ 5.2 Results and Analysis
Figure 4: Convergence curve of the EGA for task allocation. The algorithm rapidly reduces the best total path length in the early generations and converges to a near-optimal solution after approximately 92 generations, indicating stable performance.
@@
Fig. 5 and 6 provide compelling evidence for the value of the CMA-ES trajectory optimization stage. By intelligently reordering the sequence in which tasks are visited, CMA-ES effectively shortens the required travel distances and streamlines the routes, thereby significantly improving the overall efficiency and likely reducing the energy consumption of the planned missions compared to using only the initial feasible paths from Informed-RRT*. This improvement stems from the powerful adaptive search capability of CMA-ES. By iteratively updating the covariance matrix $C$, CMA-ES can learn the correlations among task points. For instance, in a dense task region, the algorithm learns that tasks within this area tend to be visited consecutively, thereby focusing the search toward solution spaces that generate ”clustered” visiting sequences. This adaptive mechanism, based on historical search information, enables the algorithm to efficiently explore and converge to globally optimal or near-optimal visiting orders, rather than merely performing local refinements, thus achieving a significant reduction in path length.
Figure 5: Quantitative comparison of total path length before and after CMA-ES trajectory optimization. The application of CMA-ES reduces the total path length from 61.2 km to 48.8 km, achieving a 20.3% improvement in route efficiency for this specific mission scenario.
@@
Table 1: Extensibility Analysis
@@
Figure 6: Visual comparison of a representative vehicle trajectory before and after CMA-ES optimization. The optimized trajectory (blue) is significantly shorter and more direct than the unoptimized one (red), while both paths successfully avoid obstacles (gray shaded areas) and connect the base station and assigned task points.
@@
Figure 7: Final cooperative mission plan for the heterogeneous UAV-UGV fleet in the simulated disaster environment. The trajectories of UAVs (solid lines) and UGVs (dashed lines) originate from the central base, visit their assigned task locations, and return to the base, while navigating around obstacle regions. This demonstrates the framework’s ability to generate a complete, feasible, and coordinated plan.
@@
5.2.1 Discussion on Results and System Limitations
The simulation results presented in Figs. 3–7 and Table 1 consistently demonstrate the superior performance of the proposed integrated framework. The EGA effectively leverages vehicle heterogeneity, assigning tasks to UAVs for rapid response and to UGVs for heavy payloads, resulting in significantly reduced mission times compared to baseline methods. The CMA-ES optimization further refines these plans, achieving notable reductions in total path length and, by extension, energy consumption.
However, a critical analysis of the results also reveals several inherent limitations of the current framework, which are important to acknowledge:
  1. 1.
    Static Environment Assumption: The framework assumes a static environment with fixed obstacles and tasks, neglecting real-world dynamics such as shifting debris, emerging tasks, or changing weather, and lacks real-time replanning capability.
  2. 2.
    Simplified Inter-Vehicle Coordination: Dynamic collision avoidance is not explicitly handled in the high-level planner, relying instead on lower-level controllers, which may lead to conflicts in dense operations.
  3. 3.
    Idealized System Models: Simulations assume perfect communication and ideal vehicle dynamics, without accounting for communication delays, packet loss, or control noise prevalent in disaster zones.
  4. 4.
    Abstracted Energy Model: Energy consumption is approximated by path length ($C(P_{i})$), omitting effects of speed, terrain, and environmental conditions on actual energy usage.
These limitations, while necessary for the tractability of the initial problem, highlight the gap between the simulated environment and the complexities of real-world deployment. They also naturally point towards the most promising directions for future research, which are discussed in the following section.
6 Conclusion
This paper has proposed a cooperative trajectory optimization and task allocation framework for heterogeneous UAV-UGV teams in post-disaster medical rescue scenarios. By integrating an EGA for global task distribution, Informed-RRT* for local collision-free path generation, and CMA-ES for sequence refinement, the framework effectively addresses the challenges of complex, obstacle-ridden environments. Simulation results demonstrate that the proposed method significantly reduces total mission completion time and overall path length compared to baseline strategies, validating its effectiveness in improving rescue efficiency. The framework successfully leverages the complementary strengths of UAVs and UGVs, enabling intelligent task assignment and generating feasible, optimized trajectories. The systematic approach—separating global allocation, path feasibility, and local sequence optimization—provides a scalable and robust solution for coordinating heterogeneous robot teams.
Despite its promising results, this work has several limitations. The assumption of a static environment limits its applicability; future work will focus on developing a dynamic replanning mechanism for real-time response to new tasks and moving obstacles.
7 Code Availability